# Universal Counterdiabatic Driving

### v1

@@ Abstract
Local counterdiabatic (CD) driving provides a systematic way of constructing a control protocol to approximately suppress the excitations resulting from changing some parameter(s) of a quantum system at a finite rate. However, designing CD protocols typically requires knowledge of the original Hamiltonian a priori. In this work, we design local CD driving protocols in Krylov space using only the characteristic local time scales of the system set by e.g. phonon frequencies in materials or Rabi frequencies in superconducting qubit arrays. Surprisingly, we find that convergence of these universal protocols is controlled by the asymptotic high frequency tails of the response functions. This finding hints at a deep connection between the long-time, low frequency response of the system controlling non-adiabatic effects, and the high-frequency response determined by the short-time operator growth and the Krylov complexity, which has been elusive until now.
@@ I Introduction
Engineering CD protocol requires implementation of the Adiabatic Gauge Potential (AGP) Sels and Polkovnikov (2017), which is generally unknown for interacting systems. Moreover for chaotic systems it is even poorly defined being a very nonlocal operator with the squared norm scaling linearly with the Hilbert space-dimension or even faster Pandey et al. (2020). The AGP is usually better defined if it only suppresses non-adiabatic transitions to/from the ground state, but even then it is generally a highly nonlocal operator, especially if the spectrum is gapless. Therefore finding the CD protocol exactly, let alone implementing it, is impractical beyond small systems. This has motivated the development of a variational approach: local counterdiabatic driving Sels and Polkovnikov (2017), where the suppression of excitations is not exact, and some error is introduced. This error can be systematically reduced by increasing the variational manifold of extra control terms. The variational principle does not rely on knowledge of the eigenstates of the instantaneous Hamiltonian it has to follow, but does need precise knowledge of this Hamiltonian. A particularly convenient variational ansatz can be obtained using operators from the so-called Krylov space Claeys et al. (2019); Bhattacharjee (2023); Takahashi and Del Campo (2024).
@@
In practice the Hamiltonian might not be known exactly due to presence of some uncontrolled terms. In this work, we ask whether one can find good “universal” CD protocols, which do not require any detailed information about the Hamiltonian of the system except for overall local energy scales, setting the natural time units. We find that it is generally possible to define such a universal CD protocol, but surprisingly its complexity is controlled not by the low-energy excitations (small gaps) as one might guess, but rather the high energy states which are artificially excited by the error in the approximate drive deviating from the true CD protocol. This result has potential implications beyond counterdiabatic driving, hinting at a connection between the short- and long-time dynamics of Hamiltonian systems. Note that while we focus on quantum systems here, a similar analysis equally applies to classical systems.
@@
The paper is organized as follows: In Section II we review CD driving and discuss how finding variational CD protocols in Krylov space can be cast as the mathematical problem of approximating $1/x$ by odd polynomials. Then we discuss how we can use this to find universal protocols, and how the high-frequency response of a system can limit the convergence of this procedure. In Section III we give examples and numerically confirm our quantitative predictions about the effect of the high frequency response on convergence. Finally, we conclude in Section IV, highlighting the main results and suggesting possible future directions of exploration.
@@ II Approximate Local CD Driving
where $A_{\lambda}$ is the so-called adiabatic gauge potential (AGP). The extra counter term in the Hamiltonian suppresses non-adiabatic transitions which lead to excitations from the ground state for any rate $\dot{\lambda}$. For a non-degenerate spectrum the matrix elements of this operator are Kolodrubetz et al. (2017)
@@
where $\omega_{mn}=E_{m}-E_{n}$ and we set $\hbar=1$. In practice $\mu$ represents the inverse time cutoff. At finite value of $\mu$ the CD driving with this AGP suppresses transitions outside of the shell $|E_{n}-E_{m}|<\mu$. In particular, if we are interested in suppressing transitions from the ground state one can set $\mu$ to be the minimal energy gap, which is usually significantly (exponentially) larger than the spectral gap in the middle of the spectrum.
@@
[…]
@@
is the spectral function. Here the probabilities $\rho_{n}$ represent some stationary statistical ensemble, for example the Gibbs distribution. If we are interested in the ground state, then we can choose $\rho_{n}=\delta_{n0}$. Another convenient choice is the infinite temperature ensemble $\rho_{n}=1/\mathcal{D}$, with $\mathcal{D}$ the Hilbert space dimension. In this case $\chi_{\lambda}$ coincides with the square of the standard Frobenius norm of the AGP. From Eq. (2) we see that the fidelity susceptibility is determined by the nearby energy states or equivalently according to Eq. (3) by the low-frequency scaling of the spectral function. In particular, for the ground state $\rho_{n}=\delta_{n0}$
@@
where $\Delta=E_{1}-E_{0}$ is the energy gap. This result is of course expected as it tells us that the complexity of the CD protocol is dictated by the minimal gap along the adiabatic path. The same gap also sets the minimal annealing time in a standard adiabatic protocol and hence determines the adiabatic complexity in time.
@@
The AGP is defined as the exact minimum of the action, where, in the limit $\mu\to 0$, $[G_{\lambda}(A_{\lambda}),H]=0$. The approximate AGP can be found by minimizing the action over some space of local accessible operators.
It has been observed that a natural variational basis in which to construct the approximate AGP is to make an expansion in Krylov space Claeys et al. (2019); Sels and Polkovnikov (2023a); Takahashi and Del Campo (2024); Bhattacharjee (2023), which corresponds to choosing a series of operators generated by the action of $[H,\boldsymbol{\cdot}\ ]$ on $\partial_{\lambda}H$ as the variational basis:
@@
The Krylov space is constructed by a proper Gram-Schmidt orthogonalization procedure in the space of nested commutators. One determines the $\alpha_{k}$ by substituting $X=A_{\lambda}^{(\ell)}$ into Eq. (5) and choosing the $\alpha_{k}$ to minimize the action. This commutator structure also has the advantage that these terms can be realized by Floquet engineering Bukov et al. (2015); Goldman and Dalibard (2014); Choi et al. (2020).
@@ II.1 Variational AGP as polynomial fitting
where $g(x)$ is the filter function satisfying $g(x\gg 1)=1$ and $g(x\to 0)\to 0$. For the regularization scheme used in Eq. (2) this function is given by $g(x)=x^{2}/(1+x^{2})$. This function is necessary to make the solution to Eq. (8) meaningful at small $\omega$, where any finite degree polynomial vanishes while $1/\omega$ diverges. One can consider other choices as well, see e.g. Ref. [Lim et al., 2024]. For example, if there is a hard ground state gap $\Delta$ it is convenient to choose $\mu=\Delta$ and $g(x)=\theta(|x|-1)$, where $\theta(x)$ is the Heaviside step function. As it becomes clear from the following discussion our analysis does not depend on the details of $g(x)$. Clearly Eq. (8) also has to be regularized in the UV limit $\omega\to\infty$ where conversely any polynomial diverges, while $1/\omega\to 0$. This regularization will control the convergence of the AGP ansatz in Krylov space.
@@
[…]
@@
Observe that in Eq. (9) all of the information about the particular system enters via the spectral function $\Phi_{\lambda}(\omega)$. For a universal CD protocol we would like to develop a variational procedure which is independent of the details of the system. To do this, we will replace the exact spectral function with the following rectangular function:
@@
Here $\mu$ is the low-frequency cutoff, which plays the same role as the cutoff in the filter function $g(\omega/\mu)$, so we keep the same notation. It determines the lowest frequency scale below which the approximate AGP does not follow the exact one. Likewise, the scale $\Omega$ plays the role of the upper frequency cutoff. Clearly, the universal variational AGP construction converges if by increasing the order of approximation $\ell$ we are able to decrease the low frequency $\mu$ while keeping $\Omega$ sufficiently large such that the exact cost function (9) remains small. The advantage of the universal approach is that the information about the system only enters through the scaling of the exact spectral function in the limits $\omega\sim\mu\to 0$ and $\omega\sim\Omega\to\infty$, which in turn determine the scaling of the error with $\ell$. The variational optimization becomes a purely mathematical problem of fitting the function $-1/\omega$ by odd polynomials in $\omega$, within the range $\mu\leq\omega\leq\Omega$. This procedure is schematically illustrated in Figure 1. The local CD protocol does not suppress transitions to the states with $|\omega_{mn}|<\mu$ and conversely induces transitions at $|\omega_{mn}|>\Omega$. After the optimal fitting in this window is achieved the parameters $\mu$ and $\Omega$ have to be optimized for each $\ell$. In this way the number of independent variational parameters is reduced from $\ell$ to only two.
@@
The expansion for the AGP in Eq. (6) in terms of nested commutators leads to the polynomial approximation of the function $1/\omega$ as it is evident from Eq. (9). It is more convenient instead to use an expansion in different polynomials. The reason is that the monomials of Eq. (7) are difficult to work with numerically due to the exponential growth of the coefficients $\alpha_{k}$. In the direct variational approach it is more convenient to work with orthogonal operators forming a Krylov space Sels and Polkovnikov (2023b); Bhattacharjee (2023); Takahashi and Del Campo (2024). It is straightforward to check that this procedure is equivalent to an approximation of $1/\omega$ in Eq. (9) in terms of polynomials which are orthogonal with respect to the spectral function $\Phi_{\lambda}(\omega)$, which depends on the details of the system. Thus it is not suitable for a universal protocol. Instead we replace Eq. (6) with a linear combination of model independent orthogonal polynomials. In this work we consider an expansion in terms of Chebyshev polynomials $T_{n}(\omega/\Omega)$. We rescale the frequency $\omega$ because the Chebyshev polynomials $T_{n}(x)$ are well-behaved and orthogonal in the interval $x\in[-1,1]$ with the weight function $1/\sqrt{1-x^{2}}$ Trefethen (2019). In the operator language this expansion corresponds to replacing the ansatz (6) with an equivalent one:
@@
where $\hat{\mathcal{L}}$ stands for the Liouvillian super-operator: ^$\hat{\mathcal{L}}O=[H,O]$. It is clear that the set of coefficients $\beta_{k}$ and $\alpha_{k}$ are uniquely related for any $\ell$. More detail about this construction is given in Appendix B.
@@
[…]
@@
An accurate polynomial approximation of $1/\omega$ in the interval $[\mu,\Omega]$ ensures that the main contributions to Eq. (9) comes from the regions $\omega\lesssim\mu$ and $\omega>\Omega$. The first, low frequency contribution is determined by the minimum gap, and more generally by the density of the low-energy states. If optimal value of $\mu(\ell)$ is a monotonically decreasing function of $\ell$, the low frequency contribution is suppressed more and more as $\ell\to\infty$. The latter, high frequency contribution is clearly determined by the high frequency tail of $\Phi_{\lambda}(\omega)$, which, in the thermodynamic limit, is typically given by
@@
The remainder of this work will be concerned with how the two parameters $\mu(\ell)$ and $\Omega(\ell)$ for the optimal protocol scale with the order of approximation $\ell$. As we shall see, not only $\Omega(\ell)$, but surprisingly also $\mu(\ell)$ depends strongly on the exponent $\alpha$: for $\alpha\leq 1$ the universal protocol does not exist in the thermodynamic limit in the sense that the UV error introduced by using higher order polynomials is too big to allow for $\mu(\ell)$ to decrease with $\ell$. Conversely for $\alpha>1$ the universal protocol works such that $\mu(\ell)\sim\ell^{1/\alpha-1}$ is a decreasing with $\ell$. The generic borderline case $\alpha=1$ requires extra care and one can achieve very slow convergence $\mu(\ell)\sim 1/\log(\ell)$ Finžgar et al. (2025). In the remainder of this paper we will explain where this result comes from and illustrate the emerging scaling of $\mu(\ell)$ and $\Omega(\ell)$ for three different situations.

### v2

@@ Abstract
Local counterdiabatic (CD) driving provides a systematic way of constructing a control protocol to approximately suppress the excitations resulting from changing some parameter(s) of a quantum system at a finite rate. However, designing CD protocols typically requires knowledge of the original Hamiltonian a priori. In this work, we design local CD driving protocols in Krylov space using only the characteristic local time scales of the system set by e.g. phonon frequencies in materials or Rabi frequencies in superconducting qubit arrays. Surprisingly, we find that convergence of these universal protocols is controlled by the asymptotic high frequency tails of the response functions. This finding hints at a deep connection between the long-time, low frequency response of the system controlling non-adiabatic effects, and the high-frequency response determined by the short-time operator growth and the Krylov complexity. We make this connection concrete by showing how, for a representative integrable model, we may extract long-time universal behaviour of the correlation functions from a short-time expansion of the dynamics using a system-independent universal protocol.
@@ I Introduction
Engineering CD protocols requires implementation of the Adiabatic Gauge Potential (AGP) Sels and Polkovnikov (2017), which is generally unknown for interacting systems. Moreover for chaotic systems it is even poorly defined being a very nonlocal operator with the squared norm scaling linearly with the Hilbert space-dimension or even faster Pandey et al. (2020a). The AGP is usually better defined if it only suppresses non-adiabatic transitions to/from the ground state, but even then it is generally a highly nonlocal operator, especially if the spectrum is gapless. Therefore finding the CD protocol exactly, let alone implementing it, is impractical beyond small systems. This has motivated the development of a variational approach: local counterdiabatic driving Sels and Polkovnikov (2017), where the suppression of excitations is not exact, and some error is introduced. This error can be systematically reduced by increasing the variational manifold of extra control terms. The variational principle does not rely on knowledge of the eigenstates of the instantaneous Hamiltonian it has to follow, but does need precise knowledge of this Hamiltonian. A particularly convenient variational ansatz can be obtained using operators from the so-called Krylov space Claeys et al. (2019); Bhattacharjee (2023); Takahashi and Del Campo (2024).
@@
In practice the Hamiltonian might not be known exactly due to presence of some uncontrolled terms. In this work, we ask whether one can find good “universal” CD protocols, which do not require any detailed information about the Hamiltonian of the system except for overall local energy scales, which set the natural time units. We find that it is generally possible to define such a universal CD protocol, but surprisingly its complexity is controlled not by the low-energy excitations (small gaps) as one might guess, but rather the high energy states which are artificially excited by the error in the approximate drive deviating from the true CD protocol. This result has potential implications beyond counterdiabatic driving, hinting at a connection between the short- and long-time dynamics of Hamiltonian systems. Note that while we focus on quantum systems here, a similar analysis equally applies to classical systems.
@@
The paper is organized as follows: In Section II we review CD driving and discuss how finding variational CD protocols in Krylov space can be cast as the mathematical problem of approximating $1/x$ by odd polynomials. Then we discuss how we can use this to find universal protocols, and how the high-frequency response of a system can limit the convergence of this procedure. In Section III we give examples and numerically confirm our quantitative predictions about the effect of the high frequency response on convergence. We make this short- and long-time dynamics connection more explicit in Section IV, where we demonstrate how the universal protocol can be combined with a short-time expansion to extract universal behaviour in long-time correlation function of physical observables, or equivalently the low frequency asymptote of their spectral functions. Finally, we conclude in Section V, highlighting the main results and suggesting possible future directions of exploration.
@@ II Approximate Local CD Driving
where $A_{\lambda}$ is the so-called adiabatic gauge potential (AGP). The extra counter term in the Hamiltonian suppresses non-adiabatic transitions which lead to excitations from the ground state for any rate $\dot{\lambda}$. For a non-degenerate spectrum the matrix elements of this operator are (see Eq. (63) of [Kolodrubetz et al., 2017])
@@
where $\omega_{mn}=E_{m}-E_{n}$, we have set $\hbar=1$, where we have inserted a regularization $\mu$ so that the matrix elements remains well defined as $\omega_{mn}\rightarrow 0$. Because this the regularization plays little role for $\omega_{mn}\gtrsim\mu$, doing CD driving with this AGP suppresses transitions between states $m\leftrightarrow n$ with $|E_{n}-E_{m}|\lesssim\mu$. Another way to think of $\mu$ is as the (inverse) time cutoff, above which all non-equal time correlation functions are truncated. In particular, if we are only interested in suppressing transitions from the ground state, we can set $\mu$ to be the gap between the ground and first excited states, which is usually significantly (exponentially) larger than the spectral gap in the middle of the spectrum. If this were realized, ground states could be prepared adiabatically with arbitrarily high fidelity.
@@
$\chi_{\lambda}=\|A_{\lambda}\|^{2}=\lim_{\mu\to 0}{1\over 4\pi}\int_{-\infty}^{\infty}d\omega{\omega^{2}\over(\omega^{2}+\mu^{2})^{2}}\Phi_{\lambda}(\omega),$ (3)
@@
is the spectral function. Here the probabilities $\rho_{n}$ represent some stationary statistical ensemble, for example the Gibbs distribution at finite temperature. If we are interested in the ground state, then we can choose $\rho_{n}=\delta_{n0}$. Another convenient choice is the infinite temperature ensemble $\rho_{n}=1/\mathcal{D}$, with $\mathcal{D}$ the Hilbert space dimension. In this case $\chi_{\lambda}$ coincides with the square of the standard Frobenius norm of the AGP. From Eq. (2) we see that the fidelity susceptibility is determined by the nearby energy states or equivalently according to Eq. (3) by the low-frequency scaling of the spectral function. In particular, for the ground state $\rho_{n}=\delta_{n0}$
@@
where $\Delta=E_{1}-E_{0}$ is the energy gap. This result tells us that the complexity of the CD protocol is dictated by the minimal gap along the adiabatic path, as expected. The same gap also sets the minimal annealing time in a standard adiabatic protocol and hence determines the adiabatic complexity in time.
@@
The AGP is defined as the exact minimum of the action, where, in the limit $\mu\to 0$, $G_{\lambda}(A_{\lambda}),H]=0$. The approximate AGP can be found by minimizing the action over some space of local accessible operators.
It has been observed that a natural variational basis in which to construct the approximate AGP is to make an expansion in Krylov space Claeys et al. (2019); Sels and Polkovnikov (2023a); Takahashi and Del Campo (2024); Bhattacharjee (2023), which corresponds to choosing a series of operators generated by the action of $H,\boldsymbol{\cdot}\ ]$ on $\partial_{\lambda}H$ as the variational basis:
@@
The Krylov space is constructed by a proper Gram-Schmidt orthogonalization procedure in the space of nested commutators. One determines the $\alpha_{k}$ by substituting $X=A_{\lambda}^{(\ell)}$ into Eq. (5) and choosing the $\alpha_{k}$ to minimize the action. Although these commutators may appear to be out of reach in a realistic experimental setting, they may be realized via the techniques of Floquet engineering Bukov et al. (2015); Goldman and Dalibard (2014); Choi et al. (2020) in which one does not need to couple to any terms not present in the Hamiltonian. Instead, rapid oscillation of the original couplings at high frequency can reproduce the commutators in an effective Floquet Hamiltonian.
@@ II.1 Variational AGP as polynomial fitting
where $g(x)$ is the filter function satisfying $g(x\gg 1)=1$ and $g(x\to 0)\to 0$. For the regularization scheme used in Eq. (2) this function is given by $g(x)=x^{2}/(1+x^{2})$. This function is necessary to make the solution to Eq. (8) meaningful at small $\omega$, where any finite degree polynomial vanishes while $1/\omega$ diverges. One can consider other choices as well, see e.g. Ref. [Lim et al., 2024]. For example, if there is a hard ground state gap $\Delta$ another choice of this regularization might be to choose $\mu=\Delta$ and $g(x)=\theta(|x|-1)$, where $\theta(x)$ is the Heaviside step function. As it becomes clear from the following discussion our analysis does not depend on the details of $g(x)$. Clearly Eq. (8) also has to be regularized in the UV limit $\omega\to\infty$ where conversely any polynomial diverges, while $1/\omega\to 0$. This regularization will turn out to control the convergence of the AGP ansatz in Krylov space.
@@
$\mathcal{C}_{\ell}=\int d\omega\left(\sum_{k=1}^{\ell}\alpha_{k}\omega^{2k-1}+{1\over\omega}\,g(\omega/\mu)\right)^{2}\Phi_{\lambda}(\omega).$ (9)
@@
Crucially, in Eq. (9) all of the information about the particular system enters only via the spectral function $\Phi_{\lambda}(\omega)$. For a universal CD protocol we would like to develop a variational procedure which is independent of the details of the system. To do this, we will replace the exact spectral function with the following rectangular function:
@@
Here $\mu$ is the low-frequency cutoff, which plays the same role as the cutoff in the filter function $g(\omega/\mu)$, so we keep the same notation. It determines the lowest frequency scale below which the approximate AGP does not follow the exact one. Likewise, the scale $\Omega$ plays the role of the upper frequency cutoff.
The universal variational AGP construction will converge if by increasing the order of approximation $\ell$, we are able to decrease the low frequency cutoff $\mu$ (i.e. suppress increasingly low energy excitations) while keeping $\Omega$ sufficiently large such that the exact cost function (9) remains small. The advantage of the universal approach is that the information about the system only enters through the scaling of the exact spectral function in the limits $\omega\sim\mu\to 0$ and $\omega\sim\Omega\to\infty$, which in turn determine the scaling of the error with $\ell$. The variational optimization becomes a purely mathematical problem of fitting the function $1/\omega$ by odd polynomials in $\omega$, within the range $\mu\leq\omega\leq\Omega$. This procedure is schematically illustrated in Figure 1. The local CD protocol fails to suppress transitions to the states with $|\omega_{mn}|<\mu$ and conversely induces transitions at $|\omega_{mn}|>\Omega$. After the optimal fitting in this window is achieved by e.g. least squares the parameters $\mu$ and $\Omega$ have to be optimized for each $\ell$. In this way the number of independent variational parameters is reduced from $\ell$ to only two.
@@
The expansion for the AGP in Eq. (6) in terms of nested commutators leads to the polynomial approximation of the function $1/\omega$ from Eq. (9). It is more convenient instead to use an expansion in different polynomials. The reason is that the monomials of Eq. (7) are difficult to work with numerically due to the exponential growth of the coefficients $\alpha_{k}$. In the direct variational approach it is most convenient to work with orthogonal operators forming a Krylov space Sels and Polkovnikov (2023b); Bhattacharjee (2023); Takahashi and Del Campo (2024). It is straightforward to check that this procedure is equivalent to an approximation of $1/\omega$ in Eq. (9) in terms of polynomials which are orthogonal with respect to the spectral function $\Phi_{\lambda}(\omega)$, which depends on the details of the system. Thus it is not suitable for a universal protocol. Instead we replace Eq. (6) with a linear combination of model independent orthogonal polynomials. In this work we consider an expansion in terms of Chebyshev polynomials $T_{n}(\omega/\Omega)$. We rescale the frequency $\omega$ because the Chebyshev polynomials $T_{n}(x)$ are well-behaved and orthogonal in the interval $x\in[-1,1]$ with the weight function $1/\sqrt{1-x^{2}}$ Trefethen (2019). In the operator language this expansion corresponds to replacing the ansatz (6) with an equivalent one:
@@
where $\hat{\mathcal{L}}$ stands for the Liouvillian super-operator: $\hat{\mathcal{L}}O=[H,O]$. It is clear that the set of coefficients $\beta_{k}$ and $\alpha_{k}$ are uniquely related for any $\ell$. More detail about this construction is given in Appendix B.
@@
$\tilde{\mathcal{C}}_{\ell}(\zeta)=\int_{\zeta}^{1}dx\left(\sum_{k=1}^{\ell}\beta_{k}T_{2k-1}(x)+{1\over x}\right)^{2},$ (13)
@@
An accurate polynomial approximation of $1/\omega$ in the interval $\mu,\Omega]$ ensures that the main contributions to Eq. (9) comes from the regions $\omega\lesssim\mu$ and $\omega>\Omega$. The first, low frequency contribution is determined by the minimum gap, and more generally by the density of the low-energy states. If optimal value of $\mu(\ell)$ is a monotonically decreasing function of $\ell$, the low frequency contribution is suppressed more and more as $\ell\to\infty$. The latter, high frequency contribution is clearly determined by the high frequency tail of $\Phi_{\lambda}(\omega)$, which, in the thermodynamic limit, is typically given by
@@
The remainder of this work will be concerned with how the two parameters $\mu(\ell)$ and $\Omega(\ell)$ for the optimal protocol scale with the order of approximation $\ell$. As we shall see, not only $\Omega(\ell)$, but surprisingly also $\mu(\ell)$ depends strongly on the exponent $\alpha$: for $\alpha<1$ the universal protocol does not exist in the thermodynamic limit in the sense that the UV error introduced by using higher order polynomials is too big to allow for $\mu(\ell)$ to decrease with $\ell$. Conversely for $\alpha>1$ the universal protocol works such that $\mu(\ell)\sim\ell^{1/\alpha-1}$ decreases with $\ell$. The generic borderline case $\alpha=1$ requires extra care and one can achieve very slow convergence $\mu(\ell)\sim 1/\log(\ell)$ Finžgar et al. (2025). In the remainder of this paper we will explain where this result comes from and illustrate the emerging scaling of $\mu(\ell)$ and $\Omega(\ell)$ for three different situations.